# Uncertainty-driven Embedding Convolution

### v2

@@ 3.3 Uncertainty-aware Similarity Estimation
Measuring distances between embeddings is fundamental to many downstream applications. While distributional distances like KL divergence (Kullback & Leibler, 1951) or Wasserstein distance (Chhachhi & Teng, 2023; Gelbrich, 1990) are theoretically well-founded, their computational cost is often prohibitive in large-scale settings. To address this, we propose a lightweight, principled similarity estimator.
Let the ensembled embeddings for $\mathbf{x}$ and $\mathbf{x^{\prime}}$ be $\mathbf{q}=\mathbf{z}(\mathbf{x})\sim\mathcal{N}(\bm{\mu}_{\mathbf{q}},\bm{\Sigma}_{\mathbf{q}})$ and $\mathbf{c}=\mathbf{z}(\mathbf{x^{\prime}})\sim\mathcal{N}(\bm{\mu}_{\mathbf{c}},\bm{\Sigma}_{\mathbf{c}})$, respectively. Our similarity measure $s$ is grounded in the properties of the base embedding models, which are typically trained with contrastive objectives on $\ell_{2}$-normalized outputs. This training process naturally results in mean vectors $\bm{\mu}_{\mathbf{q}}$, $\bm{\mu}_{\mathbf{c}}$ that are already near unit-norm, which in turn makes our small-variance assumption highly plausible. Along with it, we explicitly normalize the mean vectors, allowing us to approximate cosine similarity with the dot product, $s\approx\mathbf{q}^{\top}\mathbf{c}$. This operation preserves the distribution’s tractable Gaussian form, and the intrinsically small variance ensures the approximation is accurate. We then model the resulting dot product $s$ as a Gaussian via moment matching (Randone et al., 2024; Mallik & Sagias, 2011):
@@
The $\hat{s}$ incorporates predictive uncertainty in a lightweight manner without additional sampling. Beyond its computational efficiency, the estimator $\hat{s}$ is theoretically grounded, as it approximates the squared $2$-Wasserstein distance with provably bounded error, ensuring consistent ranking behavior.
@@
Hence, ranking by $\hat{s}$ is equivalent to minimizing $W_{2}^{2}$, up to $O(\varepsilon^{2})$ error.
@@ 4 Experiments
Figure 3: Performance on MIRACL Subset across ensemble methods. The oracle represents the upper bound by selecting the best language-specific model per language. UEC achieves performances comparable to the oracle and even surpasses in some cases, with particularly strong gains in AUC@10.
@@ 4.1 MIRACL Subset
We first build a toy example with language-specialized models: E5 (Wang et al., 2022) (English/Russian), BGE-zh (Xiao et al., 2023) (Chinese), and MATRYOSHKA (Kusupati et al., 2024; Henderson et al., 2017) (Arabic). From the MIRACL Hard Negative dataset (Zhang et al., 2023), we sample 25 query-passage pairs each in English, Russian, Chinese, and Arabic. Baselines include uniform and weighted ensembles, along with an oracle upper bound that selects the best model per language. We report nDCG@10 and Recall@10 for retrieval and AUC@10 for uncertainty (Enevoldsen et al., 2025). Full experimental details are in Appendix B.3.
@@ 4.2.2 Classification
We evaluate on five classification datasets covering finance, law, poetry, multilingual intent, and social media: FinancialPhrasebankClassification (Malo et al., 2014), SwissJudgementClassification (Niklaus et al., 2021), PoemSentimentClassification (Sheng & Uthus, 2020), MassiveIntentClassification (FitzGerald et al., 2022), and TweetTopicSingleClassification (Antypas et al., 2022). Accuracy, F1 score, and AUROC are reported as evaluation metrics, where AUROC (Kuhn et al., 2023) serves as a key quantitative measure to assess how well the model estimates the degree of uncertainty in distinguishing between classes.
@@ 4.4 Ablation Study
Table 5: Ablation results for UEC’s components: uncertainty-aware similarity (Unc Sim) and adaptive convolution (Unc Conv). Removing either degrades performance, with the joint removal yielding the worst results.
@@ 5 Conclusion and Discussion
We propose the Uncertainty-driven Embedding Convolution (UEC) framework, which converts pre-trained deterministic embeddings into probabilistic representations and adaptively ensembles them using estimated uncertainty, incorporating an uncertainty-aware similarity function. Experiments on diverse tasks show that UEC consistently outperforms strong baselines by leveraging complementary embedding strengths. While our study focuses on text embeddings and uses diagonal LA for uncertainty estimation, extending UEC to multimodal settings and exploring more scalable or advanced uncertainty methods remain promising future directions. Limitations include the need for embeddings of the same dimensionality and potential risks when applied to biased or incomplete data, underscoring the importance of responsible deployment with bias mitigation and transparent uncertainty communication. Overall, UEC enhances robustness and reliability in NLP systems, particularly in high-stakes domains.
@@ Appendix A Derivation and Proof
A.1 Derivation of Bayes-optimal Weights for Uncertainty-driven Embedding Convolution
In this section, we provide a detailed derivation of the optimal ensemble coefficients used in the uncertainty-driven embedding convolution described in Section 3.2. Specifically, we aim to derive a closed-form solution for the optimal coefficients $\bm{\pi}_{k}(\mathbf{x})$ that minimize an uncertainty-aware surrogate loss reflecting the expected distance between positive pairs.
@@ A.1 Derivation of Bayes-optimal Weights for Uncertainty-driven Embedding Convolution
Let there be $K$ independent embedding models. Each model $k\in\{1,\dots,K\}$ produces a probabilistic embedding $\mathbf{z}_{k}(\mathbf{x})\sim\mathcal{N}(\bm{\mu}_{k}(\mathbf{x}),\bm{\Sigma}_{k}(\mathbf{x}))$ for input $\mathbf{x}$, where $\bm{\mu}_{k}(\mathbf{x})\in\mathbb{R}^{d}$ is the mean and $\bm{\Sigma}_{k}(\mathbf{x})\in\mathbb{R}^{d\times d}$ is the covariance matrix. Similarly, we define $\mathbf{z}_{k}(\mathbf{x}^{\prime})\sim\mathcal{N}(\bm{\mu}_{k}(\mathbf{x}^{\prime}),\bm{\Sigma}_{k}(\mathbf{x}^{\prime}))$ for a positive example $\mathbf{x}^{\prime}$.
We define the ensembled embedding as a weighted combination of model-specific embeddings:
$\mathbf{z}(\mathbf{x})=\sum_{k=1}^{K}\bm{\pi}_{k}(\mathbf{x})\cdot\mathbf{z}_{k}(\mathbf{x}),\quad\text{with }\sum_{k=1}^{K}\bm{\pi}_{k}(\mathbf{x})=\mathbf{1},\quad\bm{\pi}_{k}(\mathbf{x})\geq 0.$
@@
Text embeddings are often trained with contrastive losses such as InfoNCE (Oord et al., 2018), which aim to bring semantically similar pairs closer while pushing dissimilar pairs apart. In the case of $\ell_{2}$-normalized features, the squared Euclidean distance is closely related to cosine similarity via the identity: $\|\mathbf{u}-\mathbf{v}\|^{2}=2(1-\mathbf{u}^{\top}\mathbf{v})$ (Wang & Isola, 2020). This connection reveals that minimizing squared distance effectively encourages high cosine similarity, aligning well with the objectives of contrastive learning.
Furthermore, recent theoretical analyses (Wang & Isola, 2020) show that contrastive objectives optimize two essential properties: (1) alignment, by pulling positive pairs together, and (2) uniformity, by encouraging embeddings to spread uniformly on the hypersphere. These principles are naturally supported by using the squared loss in the normalized feature space.
Leveraging this insight, we adopt the squared Euclidean loss as a principled surrogate for contrastive objectives. Unlike InfoNCE, which requires negative sampling and temperature scaling, the squared loss admits a simpler and deterministic formulation while still capturing the key inductive biases of alignment and uniformity. For a positive pair $(\mathbf{x},\mathbf{x}^{\prime})$, and for each embedding model $k$, we define the surrogate loss as the expected squared distance between probabilistic embeddings:
$\mathcal{L}_{\text{sur}}(\bm{\pi};\mathbf{x},\mathbf{x}^{\prime})=\sum_{k=1}^{K}\bm{\pi}_{k}(\mathbf{x})\cdot\mathbb{E}_{\mathbf{z}_{k}(\mathbf{x})}\left[\|\mathbf{z}_{k}(\mathbf{x})-\mathbf{z}_{k}(\mathbf{x}^{\prime})\|^{2}\right].$
Expected Distance Between Gaussians
To compute the surrogate loss, we evaluate the expected squared distance between two independent Gaussian embeddings:
$\mathbb{E}[\|\mathbf{z}_{k}(\mathbf{x})-\mathbf{z}_{k}(\mathbf{x}^{\prime})\|^{2}],$
where $\mathbf{z}_{k}(\mathbf{x})\sim\mathcal{N}(\bm{\mu}_{k}(\mathbf{x}),\bm{\Sigma}_{k}(\mathbf{x}))$ and $\mathbf{z}_{k}(\mathbf{x}^{\prime})\sim\mathcal{N}(\bm{\mu}_{k}(\mathbf{x}^{\prime}),\bm{\Sigma}_{k}(\mathbf{x}^{\prime}))$.
Using the identity $\|\mathbf{a}-\mathbf{b}\|^{2}=\|\mathbf{a}\|^{2}+\|\mathbf{b}\|^{2}-2\mathbf{a}^{\top}\mathbf{b}$, the expectation becomes:
$\mathbb{E}[\|\mathbf{z}_{k}(\mathbf{x})-\mathbf{z}_{k}(\mathbf{x}^{\prime})\|^{2}]$ $=\mathbb{E}[\|\mathbf{z}_{k}(\mathbf{x})\|^{2}]+\mathbb{E}[\|\mathbf{z}_{k}(\mathbf{x}^{\prime})\|^{2}]-2\mathbb{E}[\mathbf{z}_{k}(\mathbf{x})^{\top}\mathbf{z}_{k}(\mathbf{x}^{\prime})]$
$=\text{tr}(\bm{\Sigma}_{k}(\mathbf{x}))+\|\bm{\mu}_{k}(\mathbf{x})\|^{2}+\text{tr}(\bm{\Sigma}_{k}(\mathbf{x}^{\prime}))+\|\bm{\mu}_{k}(\mathbf{x}^{\prime})\|^{2}-2\bm{\mu}_{k}(\mathbf{x})^{\top}\bm{\mu}_{k}(\mathbf{x}^{\prime}),$
where we used two standard results for Gaussian vectors:
  • $\mathbb{E}[\|\mathbf{z}\|^{2}]=\text{tr}(\bm{\Sigma})+\|\bm{\mu}\|^{2}$, and
  • $\mathbb{E}[\mathbf{z}^{\top}\mathbf{z}^{\prime}]=\mu^{\top}\mu^{\prime}$ for independent $\mathbf{z}\sim\mathcal{N}(\bm{\mu},\bm{\Sigma}),\ \mathbf{z}^{\prime}\sim\mathcal{N}(\bm{\mu}^{\prime},\bm{\Sigma}^{\prime})$.
Combining these yields:
$\mathbb{E}[\|\mathbf{z}_{k}(\mathbf{x})-\mathbf{z}_{k}(\mathbf{x}^{\prime})\|^{2}]=\|\bm{\mu}_{k}(\mathbf{x})-\bm{\mu}_{k}(\mathbf{x}^{\prime})\|^{2}+\text{tr}(\bm{\Sigma}_{k}(\mathbf{x}))+\text{tr}(\bm{\Sigma}_{k}(\mathbf{x}^{\prime})).$
This decomposition reveals two key components:
  • The fidelity term $\|\bm{\mu}_{k}(\mathbf{x})-\bm{\mu}_{k}(\mathbf{x}^{\prime})\|^{2}$, which measures semantic alignment between the means of the embeddings.
  • The uncertainty term $\text{tr}(\bm{\Sigma}_{k}(\mathbf{x}))+\text{tr}(\bm{\Sigma}_{k}(\mathbf{x}^{\prime}))$, which reflects how uncertain the model is about each input.
This result plays a central role in our surrogate loss formulation, as it enables UEC to quantify both how close embeddings are (fidelity) and how reliable they are (uncertainty), using closed-form computations under the Gaussian assumption.
Tractable Approximation
Directly evaluating the full surrogate loss requires computing both the fidelity term $\|\bm{\mu}_{k}(\mathbf{x})-\bm{\mu}_{k}(\mathbf{x}^{\prime})\|^{2}$ and the uncertainty terms $\operatorname{tr}(\bm{\Sigma}_{k}(\mathbf{x}))$ and $\operatorname{tr}(\bm{\Sigma}_{k}(\mathbf{x}^{\prime}))$ for every positive pair $(\mathbf{x},\mathbf{x}^{\prime})$. However, in many practical settings, this becomes computationally prohibitive, as document embeddings (corresponding to $\mathbf{x}^{\prime}$) are typically precomputed and stored in advance. This design choice is essential for scalable retrieval systems, which need to index millions of documents offline and later perform fast matching against online queries $\mathbf{x}$ at inference time.
Consequently, the embedding $\bm{\mu}_{k}(\mathbf{x}^{\prime})$ and its associated uncertainty $\bm{\Sigma}_{k}(\mathbf{x}^{\prime})$ are fixed and query-independent at runtime. Moreover, the query $\mathbf{x}$ is not known during document encoding, so it is infeasible to compute any query-dependent quantities during offline document embedding. To ensure consistency with this retrieval pipeline, we simplify the surrogate loss by retaining only the components that are computable solely from the input $\mathbf{x}$ (i.e., the query):
@@
$\min_{\pi}\sum_{k=1}^{K}\bm{\pi}_{k}(\mathbf{x})\left(\operatorname{tr}(\bm{\Sigma}_{k}(\mathbf{x}))+\|\bm{\mu}_{k}(\mathbf{x})\|^{2}\right)\quad\text{s.t.}\;\sum_{k=1}^{K}\bm{\pi}_{k}(\mathbf{x})=\mathbf{1},\;\bm{\pi}_{k}(\mathbf{x})\geq 0.$ (7)
This formulation preserves the uncertainty-awareness of our surrogate loss while adhering to the architectural constraint of precomputed document embeddings. It also allows for efficient, query-time optimization of the ensemble coefficients $\bm{\pi}_{k}(\mathbf{x})$ based only on the statistics of the query embedding.

### v3

@@ 3.3 Uncertainty-aware Similarity Estimation
Measuring distances between embeddings is fundamental to many downstream applications. While distributional distances like KL divergence (Kullback and Leibler, 1951) or Wasserstein distance (Chhachhi and Teng, 2023; Gelbrich, 1990) are theoretically well-founded, their computational cost is often prohibitive. To address this, we propose a lightweight and principled similarity estimator.
Let the ensembled embeddings for $\mathbf{x}$ and $\mathbf{x^{\prime}}$ be $\mathbf{q}=\mathbf{z}(\mathbf{x})\sim\mathcal{N}(\bm{\mu}_{\mathbf{q}},\bm{\Sigma}_{\mathbf{q}})$ and $\mathbf{c}=\mathbf{z}(\mathbf{x^{\prime}})\sim\mathcal{N}(\bm{\mu}_{\mathbf{c}},\bm{\Sigma}_{\mathbf{c}})$, respectively. Our similarity measure $s$ follows the properties of the base embedding models, which are typically trained with contrastive objectives on $\ell_{2}$-normalized outputs. This training results in mean vectors $\bm{\mu}_{\mathbf{q}},\bm{\mu}_{\mathbf{c}}$ that are near unit-norm, making the small-variance assumption plausible. We explicitly normalize the mean vectors, allowing us to approximate cosine similarity by the dot product, $s\approx\mathbf{q}^{\top}\mathbf{c}$. This preserves the Gaussian form, and the small variance ensures the approximation remains accurate. We then model the resulting dot product $s$ as a Gaussian via moment matching (Randone et al., 2024; Mallik and Sagias, 2011):
@@
The $\hat{s}$ incorporates predictive uncertainty in a lightweight manner without additional sampling. Beyond its computational efficiency, the estimator $\hat{s}$ is theoretically grounded, as it approximates the squared $2$-Wasserstein distance with provably bounded error, ensuring consistent ranking behavior. The following result holds under a mild small-variance scaling assumption, formalized in Appendix A.4.
@@
Hence, ranking by $\hat{s}$ induces the same ranking as minimizing $W_{2}^{2}$, up to $O(\varepsilon^{2})$ error.
@@ 4 Experiments
Figure 3: Performance on MIRACL Subset across ensemble methods. The oracle represents the upper bound by selecting the best language-specific model per language. UEC achieves performance comparable to the oracle and even surpasses in some cases, with particularly strong gains in AUC@10.
@@ 4.1 MIRACL Subset
We first build a toy example with language-specialized models: E5 (Wang et al., 2022) (English/Russian), BGE-zh (Xiao et al., 2023) (Chinese), and MATRYOSHKA (Kusupati et al., 2022; Henderson et al., 2017) (Arabic). From the MIRACL Hard Negative dataset (Zhang et al., 2023), we sample 100 query-passage pairs each in English, Russian, Chinese, and Arabic. Baselines include uniform and weighted ensembles, along with an oracle upper bound that selects the best model per language. We report nDCG@10 and Recall@10 for retrieval and AUC@10 for uncertainty (Enevoldsen et al., 2025). Full experimental details are in Appendix B.3.
@@ 4.2.2 Classification
We evaluate on five classification datasets covering finance, law, poetry, multilingual intent, and social media: FinancialPhrasebankClassification (Malo et al., 2014), SwissJudgementClassification (Niklaus et al., 2021), PoemSentimentClassification (Sheng and Uthus, 2020), MassiveIntentClassification (FitzGerald et al., 2022), and TweetTopicSingleClassification (Antypas et al., 2022). Accuracy, F1 score, and AUROC are reported as evaluation metrics, where AUROC (Kuhn et al., 2023) serves as a key quantitative measure to assess how well the model estimates the degree of uncertainty in distinguishing between classes.
@@ Results
4.3 Uncertainty Estimation Diagnosis
To assess whether the post-hoc probabilistic embeddings and the resulting similarity variances provide meaningful uncertainty estimates, we evaluate two aspects: (1) calibration of the Laplace-based probabilistic embedding model used in UEC, and (2) whether the similarity variance $\sigma_{s}^{2}$ faithfully reflects predictive uncertainty. For all experiments, we fit LA following the same configuration with Section 4.2, and evaluate uncertainty using a binary classification proxy task on MS MARCO.
4.3.1 Calibration of Probabilistic Embeddings
Table 4: Calibration of probabilistic embeddings derived via LA. Probabilistic embeddings outperform the deterministic ones consistently.
ECE $\downarrow$ Deter. Prob. E5 0.063 0.036 BGE 0.067 0.042 GTE 0.077 0.043 UEC 0.075 0.032
We first measure how well the Laplace-based probabilistic embeddings capture epistemic uncertainty at the model level. Across all models, LA consistently improves calibration, and UEC yields the best-calibrated predictions. These findings demonstrate that the post-hoc probabilistic embedding model provides a stable and meaningful estimate of parameter uncertainty. Further analyses—covariance structure, number of layers included in LA, and data sparsity—are provided in Appendix C, where we observe consistent robustness across all factors.
4.3.2 Variance Calibration of the Similarity Score
We next assess whether UEC’s similarity variance $\sigma_{s}^{2}$ meaningfully reflects predictive uncertainty.
Table 5: Variance calibration on SemRel24STS. UEC achieves the lowest Var-ECE, indicating that $\sigma_{s}^{2}$ faithful estimate of similarity uncertainty.
Model Var-ECE $\downarrow$ E5 0.035 BGE 0.051 GTE 0.038 UEC 0.028
Var-ECE
Since standard ECE does not apply to continuous similarity values, we introduce a variance–error metric (Var-ECE) that measures the alignment between predicted variance and empirical squared error. For each example, higher predicted variance should correspond to a larger squared residual. We group examples into variance-quantile bins and compute the absolute discrepancy between predicted variance and empirical squared error. The full derivation and extended diagnostics remain in Appendix C.2.
Results
On the multilingual SemRel24STS benchmark, UEC achieves the lowest Var-ECE (Table 5), showing that $\sigma_{s}^{2}$ provides a numerically faithful estimate of similarity uncertainty.
@@ 4.5 Ablation Study
Table 7: Ablation for UEC’s components: uncertainty-aware similarity (Unc Sim) and adaptive convolution (Unc Conv). Removing either degrades performance, with the joint removal yielding the worst results.
@@ 5 Conclusion and Discussion
We propose the Uncertainty-driven Embedding Convolution (UEC) framework, which converts pre-trained deterministic embeddings into probabilistic representations and adaptively ensembles them via Bayesian coefficient estimation and an uncertainty-aware similarity function. To the best of our knowledge, this introduces the first fully post-hoc, uncertainty-calibrated ensemble approach for embedding models, yielding consistent and reliable gains across diverse tasks. While UEC currently models only epistemic uncertainty and relies on diagonal LA, extending it to capture aleatoric or full predictive uncertainty and exploring richer posterior structures with efficient computation represent promising future directions for further improving robustness. Furthermore, UEC assumes that all embeddings share the same dimensionality and may inherit biases from the underlying models; relaxing the dimensionality constraint and developing bias-aware or fairness-adjusted uncertainty mechanisms offer important avenues for responsible real-world deployment. Finally, applying UEC to multimodal settings—where heterogeneous uncertainty sources interact across vision, speech, and text—presents an exciting opportunity to further broaden the framework’s impact.
Acknowledgement
This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT)(RS-2024-00457216).
@@ Appendix A Derivation and Proof
A.1 Detailed Derivation of the Laplace Approximation
Let $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$ denote the training set and $\mathbf{W}^{(L)}$ the parameters of the final linear layer of the embedding model. The posterior over $\mathbf{W}^{(L)}$ is given by Bayes’ rule:
$p(\mathbf{W}^{(L)}\mid\mathcal{D})=\frac{p(\mathcal{D}\mid\mathbf{W}^{(L)})p(\mathbf{W}^{(L)})}{p(\mathcal{D})}.$ (7)
The exact posterior is typically intractable. The Laplace Approximation (LA) expands the negative log-posterior around its maximizer (the MAP estimate)
$\widehat{\mathbf{W}}^{(L)}=\arg\max_{\mathbf{W}^{(L)}}p(\mathbf{W}^{(L)}\mid\mathcal{D}).$
Quadratic Expansion of the Log-Posterior
Define the log-posterior
$\ell(\mathbf{W}^{(L)})=\log p(\mathcal{D}\mid\mathbf{W}^{(L)})+\log p(\mathbf{W}^{(L)}).$
A second-order Taylor expansion around $\widehat{\mathbf{W}}^{(L)}$ gives
$\ell(\mathbf{W}^{(L)})\approx\ell(\widehat{\mathbf{W}}^{(L)})+\frac{1}{2}(\mathbf{W}^{(L)}-\widehat{\mathbf{W}}^{(L)})^{\top}\nabla^{2}\ell(\widehat{\mathbf{W}}^{(L)})(\mathbf{W}^{(L)}-\widehat{\mathbf{W}}^{(L)}).$ (8)
Since $\widehat{\mathbf{W}}^{(L)}$ is a maximizer,
$\nabla\ell(\widehat{\mathbf{W}}^{(L)})=\mathbf{0}.$
Define the negative Hessian (precision matrix)
$\mathbf{H}_{\widehat{\mathbf{W}}^{(L)}}=-\nabla^{2}\ell(\widehat{\mathbf{W}}^{(L)})=\nabla^{2}\!\big[-\log p(\mathcal{D}\mid\mathbf{W}^{(L)})-\log p(\mathbf{W}^{(L)})\big]\Big|_{\mathbf{W}^{(L)}=\widehat{\mathbf{W}}^{(L)}}.$
Thus the posterior is approximated by
$p(\mathbf{W}^{(L)}\mid\mathcal{D})\approx\mathcal{N}\!\left(\widehat{\mathbf{W}}^{(L)},\mathbf{H}_{\widehat{\mathbf{W}}^{(L)}}^{-1}\right).$
Posterior over Embeddings
Following Section 3.1, let $f(\mathbf{x})$ be an embedding model with last-layer parameters $\mathbf{W}^{(L)}$ and penultimate representation $\mathbf{h}^{(L-1)}(\mathbf{x})$. The embedding is given by the final linear projection
$\mathbf{z}(\mathbf{x})=\mathbf{W}^{(L)}\mathbf{h}^{(L-1)}(\mathbf{x}).$
Under the Gaussian posterior
$\mathbf{W}^{(L)}\sim\mathcal{N}\!\left(\widehat{\mathbf{W}}^{(L)},\mathbf{H}_{\widehat{\mathbf{W}}^{(L)}}^{-1}\right),$
the embedding $\mathbf{z}(\mathbf{x})$ becomes a Gaussian random variable since it is a linear transformation of $\mathbf{W}^{(L)}$.
Using standard results on affine transformations of Gaussian variables, the predictive distribution of the embedding is given by
$\mathbf{z}(\mathbf{x})\sim\mathcal{N}\!\left(\widehat{\mathbf{W}}^{(L)}\mathbf{h}^{(L-1)}(\mathbf{x}),\;\mathbf{h}^{(L-1)}(\mathbf{x})^{\top}\mathbf{H}_{\widehat{\mathbf{W}}^{(L)}}^{-1}\mathbf{h}^{(L-1)}(\mathbf{x})\right).$ (9)
Therefore, the deterministic embedding model is converted into a probabilistic one that outputs a Gaussian embedding for each input $\mathbf{x}$, where the mean corresponds to the original embedding and the covariance captures epistemic uncertainty induced by the last-layer parameter posterior.
Diagonal Precision Approximation
In practice, computing the full Hessian is computationally expensive. We therefore adopt a diagonal approximation of the precision matrix:
$\mathbf{H}_{\widehat{\mathbf{W}}^{(L)}}\approx\mathrm{diag}\!\left(\mathbf{H}_{\widehat{\mathbf{W}}^{(L)}}\right).$
This approximation preserves tractability while retaining model-specific epistemic uncertainty information, yielding a closed-form Gaussian embedding distribution used in the UEC framework.
A.2 Derivation of Uncertainty-driven Coefficient for Embedding Convolution
In this section, we provide a detailed derivation of the uncertainty-driven coefficients used in the embedding convolution described in Section 3.2. In contrast to a hard selection induced by a linear objective, we derive a temperature-controlled, closed-form solution by introducing entropy regularization, which yields stable and query-adaptive coefficients.
@@ A.2 Derivation of Uncertainty-driven Coefficient for Embedding Convolution
Let there be $K$ independent embedding models. Each model $k\in\{1,\dots,K\}$ produces a probabilistic embedding
$\mathbf{z}_{k}(\mathbf{x})\sim\mathcal{N}(\bm{\mu}_{k}(\mathbf{x}),\bm{\Sigma}_{k}(\mathbf{x}))$
for an input $\mathbf{x}$. Similarly, for a positive example $\mathbf{x}^{\prime}$, we define
$\mathbf{z}_{k}(\mathbf{x}^{\prime})\sim\mathcal{N}(\bm{\mu}_{k}(\mathbf{x}^{\prime}),\bm{\Sigma}_{k}(\mathbf{x}^{\prime})).$
The ensembled embedding is defined as a convex combination:
$\mathbf{z}(\mathbf{x})=\sum_{k=1}^{K}\bm{\pi}_{k}(\mathbf{x})\,\mathbf{z}_{k}(\mathbf{x}),\quad\sum_{k=1}^{K}\bm{\pi}_{k}(\mathbf{x})=1,\quad\bm{\pi}_{k}(\mathbf{x})\geq 0.$
@@
Following prior analyses of contrastive representation learning (Wang and Isola, 2020), we adopt the squared Euclidean distance as a surrogate objective. For $\ell_{2}$-normalized features, minimizing squared distance is equivalent to maximizing cosine similarity, which captures the alignment property of contrastive learning while avoiding explicit negative sampling.
For a positive pair $(\mathbf{x},\mathbf{x}^{\prime})$, we define the uncertainty-aware surrogate loss as
$\mathcal{L}_{\mathrm{sur}}(\bm{\pi};\mathbf{x},\mathbf{x}^{\prime})=\sum_{k=1}^{K}\bm{\pi}_{k}(\mathbf{x})\,\mathbb{E}\!\left[\|\mathbf{z}_{k}(\mathbf{x})-\mathbf{z}_{k}(\mathbf{x}^{\prime})\|^{2}\right].$
Expected Distance Between Gaussian Embeddings
For independent Gaussian variables $\mathbf{z}_{k}(\mathbf{x})\sim\mathcal{N}(\bm{\mu}_{k}(\mathbf{x}),\bm{\Sigma}_{k}(\mathbf{x}))$ and $\mathbf{z}_{k}(\mathbf{x}^{\prime})\sim\mathcal{N}(\bm{\mu}_{k}(\mathbf{x}^{\prime}),\bm{\Sigma}_{k}(\mathbf{x}^{\prime}))$, the expected squared distance admits the closed-form expression:
$\mathbb{E}\!\left[\|\mathbf{z}_{k}(\mathbf{x})-\mathbf{z}_{k}(\mathbf{x}^{\prime})\|^{2}\right]=\|\bm{\mu}_{k}(\mathbf{x})-\bm{\mu}_{k}(\mathbf{x}^{\prime})\|^{2}+\operatorname{tr}(\bm{\Sigma}_{k}(\mathbf{x}))+\operatorname{tr}(\bm{\Sigma}_{k}(\mathbf{x}^{\prime})).$
This decomposition separates a fidelity term, capturing semantic alignment, and an epistemic uncertainty term, reflecting model confidence.
Retrieval-feasible Approximation
In large-scale retrieval systems, document embeddings corresponding to $\mathbf{x}^{\prime}$ are precomputed and fixed, while the query $\mathbf{x}$ arrives only at inference time. To respect this constraint, we discard all document-dependent terms and retain only the query-dependent components, yielding the following query-time objective:
@@
$\min_{\bm{\pi}}\sum_{k=1}^{K}\bm{\pi}_{k}(\mathbf{x})\Big(\operatorname{tr}(\bm{\Sigma}_{k}(\mathbf{x}))+\|\bm{\mu}_{k}(\mathbf{x})\|^{2}\Big)\;-\;T\,\mathrm{H}\!\left(\bm{\pi}(\mathbf{x})\right),\quad\text{s.t. }\sum_{k}\bm{\pi}_{k}(\mathbf{x})=1,\;\bm{\pi}_{k}(\mathbf{x})\geq 0,$ (10)
where $H(\bm{\pi}(\mathbf{x})):=-\sum_{k}\bm{\pi}_{k}(\mathbf{x})\log\bm{\pi}_{k}(\mathbf{x})$ denotes the Shannon entropy.
The entropy term prevents degenerate solutions that collapse onto a single model and can be interpreted as a KL divergence to a uniform prior over models, thereby promoting robust aggregation under noisy, query-specific uncertainty.